A fund manager with the Ariri Fund Management, is contemplating between these bonds to be added to one of the firm’s balanced funds and balanced out the risks in the fund. Among the criteria he is looking for is a bond with an A and above rating, has less than 5 years to maturity and less volatile towards the changes in the interest rate. Listed are the corporate bonds available today.


BOND RATING COUPON (%) TERM TO MATURITY PAR VALUE ($)

A A 5 2 1,000

B A 7 10 1,000

C AA 4 3 1,000

D BB 6 4 1,000


The pandemic has eroded the value of money and lower down the purchasing power. He

expects that the interest rate is rising to 7 percent soon from the existing rate of 5 percent.

Based on the information given, his criteria, and modified duration; assist him to choose

one bond from the listed bonds.


1
Expert's answer
2022-01-06T10:44:25-0500

bonds with an A and above rating that have less than 5 years to maturity: A and C


Present value of bond:

"PV= C* \frac{1-(1+r)^{-t}}{r} + \frac{F}{(1+r)^t}"


for bond A:

"PV=50\cdot\frac{1-(1+0.05)^{-2}}{0.05}+\frac{1000}{(1+0.05)^2}=1000"


for bond C:

"PV=40\cdot\frac{1-(1+0.05)^{-3}}{0.05}+\frac{1000}{(1+0.05)^{3}}=972.77"


so, fund manager should choose bond A (with maximal PV)